OpenStudy (rissyroo118):

help simplify please (4^3)5 A)4^-2 B)4^3 C)4^8 D)4^15

OpenStudy (anonymous):

D

Nnesha (nnesha):

\[\huge\rm (x^n)^m = x^{n \times m}\] multiply outside exponent by inside exponent of parentheses

OpenStudy (rissyroo118):

@Nnesha i got B?

OpenStudy (anonymous):

its D 3*5=15

Nnesha (nnesha):

exponent which is inside the parentheses is 3 and outside exponent is 5  letz say n represent 3 and m represent 5 so multiply n times m = ?